# Adaptive Rank, Reduced Forgetting: Knowledge Retention in Continual Learning Vision-Language Models with Dynamic Rank-Selective LoRA

### v5

@@ B.1 Effects of Additional Adaptive Prediction Head on Visual Representations
The related continual learning methods focus on updating the representation with the backbone model (e.g., [53, 66, 61], and CoDyRA) or newly added representation alignment or prediction head [37, 58]. These two types of methods can be seen as orthogonal. We study how the representation updated by the proposed can work with the additional adaptive prediction head, relying on the aggression-based adapters in [58].
@@
We provide a more detailed analysis of the effects of the additional prediction head on the visual representations in X-TAIL setting in Table 5. Note that the Dual version of RAIL [58] stores all visual representations of all samples for all seen classes and domains in memory, it is not included in Tables 1 and 2 for a fair comparison.
@@
B.2 Analysis of Changes/Interference w.r.t. Pre-Trained Weights
In Sec. 4.3 of the main paper, Fig. 6 and Fig. 6 visualize and analyze the dynamic ranks assigned to each module. Here, we extend the analysis by comparing parameter changes to pre-trained weights using the amplification factor as in [21]. The amplification factor is given by:
@@
A higher amplification factor indicates that low-rank parameter updates amplify directions less emphasized in the pre-trained weights [21]. We visualize the amplification factors for both our method and vanilla LoRA trained on the Aircraft, EuroSAT and Oxford Pets datasets in Fig. 9.
@@
B.3 More Visualizations and Analyses on Rank allocation results of different datasets
@@
B.4 Evaluation with Other VLMs
To evaluate the broader applicability of our method, we conducted preliminary experiments with BLIP [28] on X-TAIL (Table 6). Our method consistently outperformed prior works and exceeded the pre-trained model’s upper-bound performance, demonstrating its general effectiveness across different VLMs.
@@ B.4 Evaluation with Other VLMs
B.5 Evaluation on Standard Parameter-Efficient Fine-Tuning (PEFT) Task
@@
Broader Impacts
CoDyRA’s sparse, adaptive, importance-aware low-rank subspace learning significantly reduces compute and memory demands for continual learning, enabling on-device personalization in domains such as mobile health and education and empowering resource-constrained users and organizations. As a fundamental research effort, CodyRA does not address potential biases or hazards in training data, and its incremental updates could introduce privacy risks or stealthy backdoors. We therefore recommend transparent update logging, regular audits, and differential-privacy safeguards. Although more efficient than full fine-tuning, large-scale deployment still incurs energy costs; developers should monitor resource use, leverage low-power hardware, and explore carbon offsets.

### v6

@@ B.2 Effects of Additional Adaptive Prediction Head on Visual Representations
The related continual learning methods focus on updating the representation with the backbone model (e.g., (Wang et al., 2022e; Zheng et al., 2023; Yu et al., 2024), and CoDyRA) or newly added representation alignment or prediction head (McDonnell et al., 2024; Xu et al., 2024). These two types of methods can be seen as orthogonal. We study how the representation updated by the proposed can work with the additional adaptive prediction head, relying on the aggression-based adapters in (Xu et al., 2024).
@@
We provide a more detailed analysis of the effects of the additional prediction head on the visual representations in X-TAIL setting in Table 8. Note that the Dual version of RAIL (Xu et al., 2024) stores all visual representations of all samples for all seen classes and domains in memory, it is not included in Tables 1 and 3 for a fair comparison.
@@
B.3 Analysis of Changes/Interference w.r.t. Pre-Trained Weights
In Sec. 4.3 of the main paper, Fig. 6 and Fig. 6 visualize and analyze the dynamic ranks assigned to each module. Here, we extend the analysis by comparing parameter changes to pre-trained weights using the amplification factor as in (Hu et al., 2021). The amplification factor is given by:
@@
A higher amplification factor indicates that low-rank parameter updates amplify directions less emphasized in the pre-trained weights (Hu et al., 2021). We visualize the amplification factors for both our method and vanilla LoRA trained on the Aircraft, EuroSAT and Oxford Pets datasets in Fig. 9.
@@
B.4 More Visualizations and Analyses on Rank allocation results of different datasets
@@
B.5 Evaluation with Other VLMs
To evaluate the broader applicability of our method, we conducted preliminary experiments with BLIP (Li et al., 2022) on X-TAIL (Table 9). Our method consistently outperformed prior works and exceeded the pre-trained model’s upper-bound performance, demonstrating its general effectiveness across different VLMs.
@@ B.5 Evaluation with Other VLMs
B.6 Evaluation on Standard Parameter-Efficient Fine-Tuning (PEFT) Task